# Stability of Thin Shell and Wormhole Configurations: Schwarzschild, Schwarzschild - (Anti-) de Sitter, and FLRW Spacetimes

### v1

@@ I Introduction
The structure of the paper is described as follows. In Section II, we begin by establishing and summarizing the thin shell formalism, using the equation of motion of the radius of the throat to derive stability conditions. In Section III, the stability conditions are applied to wormhole and black hole junctions composed of Schwarzschild, Schwarzschild - de Sitter, and Schwarzschild anti-de Sitter spacetimes. A taxonomy of major categories is defined, and mathematical conditions related to the geometry of the stability regions for each category are derived (this portion will build and expand [6]. Next, in Section IV we consider the construction of wormholes and black holes using the Friedmann–Lemaître–Robertson–Walker (FLRW) metric and a Schwarzschild or Schwarzschild - (anti-) de Sitter metric. We derive the extrinsic curvature and use thin shell formalism to give and categorize the stability conditions of these junctions. The taxonomic conditions from the earlier section are generalized to apply to this latter section. In Section V, we present our analysis with plots and discussion. And finally, in Section VI a summary and concluding remarks are provided.
@@ II Formalism
(a) Black hole
(b) Wormhole
Figure 1: Schwarzschild – Schwarzschild Junction: For the black hole, $M\neq 0$ everywhere and there is no stability flip. Stability region is always underneath the surface, partially intersecting the Causal Region (i.e., $10<P^{\prime}/\sigma^{\prime}<$ 1). For the wormhole, there is an asymptote and stability region is above surface A and below surface B.
(a) Black hole
(b) Wormhole
Figure 2: Schwarzschild - de Sitter – Schwarzschild Junction: Plot has been expanded out to show horizons. The de Sitter Horizon is visible, as expected for a junction containing a $\Lambda>0$  [^2] The horizon occurs at $R/m^{+}\approx 54$ which is close to $\alpha_{\rm dS}\approx 54.77$ for $C=0.001$ – see text. For the black hole, stability is defined by the vertical surface bifurcating the plot where $M=0$. Stability regions are below surface A and above surface B. A Schwarzschild - de Sitter – Schwarzschild junction implies $[C]=-0.001$ which does not fulfill the asymptote condition for any beta. Stability regions for the wormhole are below surface B and above surface A which is too small to be visible in this plot.
(a) Black hole
(b) Wormhole
Figure 3: Schwarzschild Anti-de Sitter – Schwarzschild Junction: Black hole has an asymptote as $[C]=0.001$ which does fulfill the asymptote condition. Stability regions for black hole below A and above B. Though not shown, wormhole at high $R$ is very similar to Schwarzschild – Schwarzschild and does not possess a de Sitter Horizon. Stability regions are above A and below B.
(a) Black hole
(b) Wormhole
Figure 4: Schwarzschild – Schwarzschild - de Sitter Junction: Black hole is similar to Schwarzschild - anti-de Sitter – Schwarzschild (FIG. 3) with a similar asymptote, though in this case the de Sitter Horizon is also present. Again $[C]=0.001$ in this case, fulfilling the asymptote condition. Stability regions for black hole below A and above B. Wormhole case at large $R$ is similar to Schwarzschild - de Sitter – Schwarzschild (FIG. 2). Stability regions are above A and below B.
(a) Black hole
(b) Wormhole
Figure 5: Schwarzschild – Schwarzschild - anti-de Sitter Junction: Black hole is similar to Schwarzschild - de Sitter – Schwarzschild (figure 2) but does not de Sitter Horizon. $[C]=-0.001$ and there is no asymptote. Stability regions are above A and below B.
@@
Throughout this work, we shall define $[A]\equiv A^{+}-A^{-}$ and $\bar{A}\equiv\frac{1}{2}(A^{+}+A^{-})$ for some quantity $A$. The first Darmois condition for joining a portion of $\mathcal{M}^{+}$ to a portion of $\mathcal{M}^{-}$ is (see, e.g., [4])
$[g_{ij}]=0.$ (1)
@@
(a) Black hole
(b) Wormhole
Figure 6: Schwarzschild - de Sitter – Schwarzschild - de Sitter Junction: Wormhole case at large $R$ is similar to Schwarzschild - de Sitter – Schwarzschild (figure 2). Stability regions are above A and below B. Black hole still possesses asymptote as $[C]=0$ fulfills asymptote condition in equation (27) if there exists a $\Lambda>0$. The location and shape of the asymptote differs from other cases and occurs at much higher $R$.
@@
This boundary will cause a discontinuity in the extrinsic curvature (second fundamental form) of the union of $\mathcal{M}^{+}$ and $\mathcal{M}^{-}$. The stress-energy tensor ($S_{ij}$) for this boundary can be calculated using the Lanczos equation, which is given by [4]
@@
We will treat $S_{ij}$ analogously to a perfect 4-fluid with $S_{ij}={\rm diag}(-\sigma,P,P)$ [3]. It can be shown that the energy density is
@@
If equation (1) holds and $[K_{\theta\theta}]=0$ then we refer to $\Sigma$ as a boundary surface, if $[K_{\theta\theta}]\neq 0$ then $\Sigma$ is a thin shell.
22footnotetext: The horizon occurs at $R/m^{+}\approx 54$ which is close to $\alpha_{\rm dS}\approx 54.77$ for $C=0.001$ – see text
In Section III, we will consider a junction between two spacetimes of the form,
@@
where $w$ is determined by the direction of the normal vectors. If the vectors point in the opposite directions $w=-1$ and the junction is referred to as a wormhole. If the vectors point in the same direction $w=1$ and the junction is referred to as a black hole.
@@
For this static solution, $V(R_{0})=V^{\prime}(R_{0})=0$. A stable solution will be given by $V^{\prime\prime}(R_{0})>0$. It follows from the definition of $V(R)$ that the equilibrium condition $V^{\prime}(R)=0$ for a static solution is
$\bigg{(}\frac{M}{2R}\bigg{)}^{\prime}=-\frac{2R}{M}\bigg{(}\bigg{(}\frac{(\mu]}{R}\bigg{)}\bigg{(}\frac{[\mu]}{R}\bigg{)}^{\prime}+\bigg{(}\frac{\bar{\mu}}{R}\bigg{)}^{\prime}\bigg{)}\equiv\Gamma.$ (13)
@@
$\dot{\sigma}=-2\frac{\dot{R}}{R}(\sigma+P)+\Xi$ (16)
@@
Note that the stability regions will lie above and below the surface mapped out by the transparency condition (equation 19).
Stability within the region defined by $10\leq\frac{P^{\prime}}{\sigma^{\prime}}<$ 1 is of particular interest as it is the causal region. $P^{\prime}/\sigma^{\prime}$ corresponds to the square of a sound speed of perturbations. Thus, for physical solutions it seems natural to restrict this sound speed to real, sub-luminal values, where it is causal. The legitimacy of this assumption is discussed in the conclusion.
@@ III Schwarzschild and (Anti-) de Sitter spacetimes
We begin our stability analysis by considering Schwarzschild wormhole and black hole solutions. We will also consider the effects of a positive and negative cosmological constant corresponding to Schwarzschild - de Sitter and Schwarzschild - anti-de Sitter spacetime respectively.
@@
Throughout the rest of this paper where relevant, we will only consider cases where $m^{-}<m^{+}$, which is reflected in our axis limits of our plots $10<m^{-}/m^{+}<$ 1 [6] (in general $\mu^{-}$ and $\mu^{+}$ can be swapped without consequence to the stability regions).
@@
An asymptote between these two horizons will occur if $\Upsilon=0$ [6]. A wormhole with any combination of the above spacetimes will always contain such an asymptote. A black hole may or may not contain an asymptote based on the following.
It can be seen from plotting that for small $R$ (near the event horizon) $\Upsilon$ and $M$ both have the same sign. With the exception of Schwarzschild – Schwarzschild (FIG. 1), after a certain point (before $\Upsilon=0$ or $M=0$) if $\Upsilon$ monotonically increases, $M$ monotonically decreases or vice versa. Thus, only one of these quantities can be equal to zero for a certain junction. If $M=0$ anywhere then $\Upsilon\neq 0$ and there is no asymptote, though a stability flip will still exist at $M=0$. This lets the simpler condition $M=0$ become an indicator of the nonexistence of an asymptote.
@@
Since $R$ must be positive, if $[m]$ and $[\Lambda]$ have the same sign, the condition can never be satisfied for any $R$, and so an asymptote must exist. This gives an asymptote existence condition of
@@
For a junction with at least one $\Lambda>0$ there is an upper limit to $R$. If the $R$ value which satisfies the horizon condition (equation 25) is greater than $R_{\rm dS}$ then the condition for asymptote non-existence cannot be met and an asymptote will be present. Taking $R_{\rm dS}\approx\sqrt{3/\Lambda_{\rm max}}$ gives an asymptote existence condition of
@@
For the Schwarzschild – Schwarzschild case (FIG. 1), $M$ and $\Upsilon$ approach zero for large $R$, but never reach it. Thus, there is neither an asymptote or stability flip in this simple case.
It is also apparent from plotting that $M\Upsilon$ will monotonically decrease for black holes and monotonically increase for wormholes. This behavior dictates the location of stability regions, whether above or below the surface.
Table 1: Notable Features of Black hole Parameter Spaces.
Type Asymptote at $\Upsilon=0$ Limit as $R\rightarrow\infty$ Stability regions Stable in causal region Figure
Sch – Sch No Convergent Below Yes 1
Sch – Sch-deSit Yes Divergent Below A, Above B Yes 4
Sch – Sch-Anti deSit No Convergent Below A, Above B Yes 5
Sch-deSit – Sch No Divergent Below A, Above B Yes 2
Sch-deSit – Sch-deSit Yes Divergent Below A, Above B Yes 6
Sch-deSit – Sch-Anti deSit No Divergent Below A, Above B Yes -
Sch-Anti deSit – Sch Yes Convergent Below A, Above B Yes 3
Sch -Anti deSit – Sch-deSit Yes Divergent Below A, Above B Yes -
Sch-Anti deSit – Sch-Anti deSit No Convergent Below Yes 7
FLRW k=0 – Sch No Convergent Below A, Above B Yes 8
FLRW k=0 – Sch-deSit No Convergent Below A, Above B Yes 9
FLRW k=0 – Sch-Anti deSit No Convergent Below A, Above B Yes 10
FLRW k=+1 – Sch No Divergent Below A, Above B Yes 11
FLRW k=-1 – Sch No Convergent Below A, Above B Yes 12
(a) Black hole
(b) Wormhole
Figure 7: Schwarzschild - anti-de Sitter – Schwarzschild - anti-de Sitter Junction: Both wormhole and black hole are very similar to Schwarzschild – Schwarzschild junction (figure 1). The wormhole parameter surface takes the usual form (stability above A and below B). The black hole has no asymtote and approaches 0 as $R\rightarrow\infty$. $M\neq 0$ for any $\beta$ and there is no stability flip. Stability region is always below the surface.
Table 2: Notable Features of Wormhole parameter Spaces.
Type Asymptote at $\Upsilon=0$ Limit as $R\rightarrow\infty$ Stability Regions Stable in Causal Region Figure
Sch – Sch Yes Convergent Above A, Below B No 1
Sch – Sch-deSit Yes Divergent Above A, Below B No 4
Sch – Sch-Anti deSit Yes Convergent Above A, Below B No 5
Sch-deSit – Sch Yes Divergent Above A, Below B No 4
Sch-deSit – Sch-deSit Yes Divergent Above A, Below B No 6
Sch-deSit – Sch-Anti deSit Yes Divergent Above A, Below B No -
Sch-Anti deSit – Sch Yes Convergent Above A, Below B No 3
Sch -Anti deSit – Sch-deSit Yes Divergent Above A, Below B No -
Sch-Anti deSit – Sch-Anti deSit Yes Convergent Above A, Below B No 7
FLRW k=0 – Sch Yes Convergent Above A, Below B No 8
FLRW k=0 – Sch-deSit Yes Convergent Above A, Below B No 9
FLRW k=0 – Sch-Anti deSit Yes Convergent Above A, Below B No 10
FLRW k=+1—FLRW Yes Divergent Above A, Below B No 11
FLRW k=-1—FLRW Yes Convergent Above A, Below B No 12
@@ IV FLRW spacetime
Unlike the previous stationary spacetimes, the FLRW spacetime is expanding. Here we follow the precedent established in the Swiss Cheese Cosmological Model, which considers spherical regions of Schwarzschild spacetime matched to an FLRW background [31]. In these cases, it is typical to define a hypersurface that is expanding at the same rate as the FLRW so that in the FLRW frame it is stationary apart from the evolution of the hypersurface radius defined in equation (11). The surface in the FLRW frame will be defined by $r_{\rm f}=R(\tau)$. The static solution solved for here will be one for which $\dot{R}=0$.
@@
$\Xi=\pm\Bigg{[}\dot{r}(\rho+p){\bigg{(}\bigg{|}g^{\alpha\beta}\frac{\partial F}{\partial x^{\alpha}}\frac{\partial F}{\partial x^{\beta}}\bigg{|}\bigg{)}^{-1/2}}\Bigg{]}$ (30)
@@
Here we will let $M=-[K_{\theta\theta}]=4\pi\chi^{2}\sigma$. Using this, the continuity equation can finally be expressed as
@@
(a) Black hole
(b) Wormhole
Figure 8: FLRW k=0 – Schwarzschild Junction: The black hole and wormhole graphs are similar to previous examples though does not posses a de Sitter Horizon. It is worth noting that the radius of the event horizon decreases rapidly at low $a$ due to a high value of $H^{2}$. This is much easier to see in the wormhole graph. Stability regions for the black hole are once again below surface A, above surface B. For wormhole, stability regions are above surface A and below surface B.
(a) Black hole
(b) Wormhole
Figure 9: FLRW k=0 – Schwarzschild - de Sitter Junction: Here $[C]=0$ causing the $M=0$ to be linear for the black hole. Outside of this, plots are very similiar to figure (8). For black hole, stability regions are below A and above B. For wormhole stability regions are above A and below B.
@@
For FLRW, $K_{\theta\theta}^{-}$ yields a complicated form. We simplify this by taking $K_{\theta\theta}$ at the static solution $\dot{R}=\frac{dR}{dt}\frac{dt}{d\tau}=0\implies\frac{dR}{dt}=0$ if $\frac{dt}{d\tau}\neq 0$.
@@
Though, at the static solution $V=H\chi\neq 0$ generally. $V^{\prime}$ may be be chosen to be 00 for an equilibrium point. Thus, $V^{\prime\prime}>0$ will again yield stable equilibria. Combining this with the transparency condition (equation 35) and the fact that $V^{\prime}=0$ at this point, gives stability conditions of
@@
(a) Black hole
(b) Wormhole
Figure 10: FLRW k=0 – Schwarzschild - anti-de Sitter Junction: Very similar to figure (8). $M=0$ plane is shifted to lower $\chi$ values. Stability regions for black hole are below A, above B. Stability regions for wormhole are above A, below B.
@@
(a) Black hole
(b) Wormhole
Figure 11: FLRW k=+1 – Schwarzschild Junction: For $k\neq 0$ we choose $m=0.02$. Note that $k>0$ gives $H^{2}<0$ for some values of $a(t)$. These non-physical solutions have been removed from the plots. A horizon caused by the positive curvature exists at high $\chi$ for both black hole and wormhole, though is more visible in the black hole. This horizon follows along the plane defined by $\chi=\frac{a}{m}$. At larger $m$ values the asymptote condition can be fulfilled but commonly this occurs only when $H^{2}<0$. Stability regions are below A and above B for black hole and are above A and below B for wormhole.

### v2

@@ I Introduction
The structure of the paper is described as follows. In Section II, we begin by establishing and summarizing the thin shell formalism, using the equation of motion of the radius of the throat to derive stability conditions. In Section III, the stability conditions are applied to wormhole and black hole junctions composed of Schwarzschild, Schwarzschild - de Sitter, and Schwarzschild anti-de Sitter spacetimes. A taxonomy of major categories is defined, and mathematical conditions related to the geometry of the stability regions for each category are derived (this portion builds and expands on [8]). Next, in Section IV we consider the construction of wormholes and black holes using the Friedmann–Lemaître–Robertson–Walker (FLRW) metric and a Schwarzschild or Schwarzschild - (anti-) de Sitter metric. We derive the extrinsic curvature and use the thin shell formalism to give and categorize the stability conditions of these junctions. The taxonomic conditions from the earlier section are generalized to apply to this latter section. In Section V, we present our analysis with two-dimensional plots and discussion. Finally, in Section VI, a summary and concluding remarks are provided. Additionally, we display the stability regions in three-dimensional parameter space for a few configurations of both black holes and wormholes in Appendix A.
@@ II Formalism
Figure 1: Schwarzschild – Schwarzschild Junction: The areas shaded in green and blue correspond to the stability regions for black hole and wormhole, respectively. The shaded region in gray represents the causal region in both cases. This color scheme is used consistently throughout this paper. For the black hole, $M\neq 0$ anywhere and there is no stability flip. Stability partially intersects the Causal Region. For the wormhole there is an asymptote present and stability is separated into two disjoint regions, neither of which intersect the Causal Region. For 3-D plot, see figure (13).
Figure 2: Schwarzschild - de Sitter – Schwarzschild Junction: For the black hole, there are points where $M=0$ and a stability flip is present without an asymptote. This occurs at the vertical line. A Schwarzschild - de Sitter – Schwarzschild junction implies [C]=0.001[C]=-0.001 which does not fulfill the asymptote condition for any beta. For both plots the effects of the de Sitter Horizon can be seen by the divergence around $\alpha\approx 54$. For 3-D plot, see figure (14)
Figure 3: Schwarzschild Anti-de Sitter – Schwarzschild Junction: Black hole has an asymptote as $[C]=0.001$ which does fulfill the asymptote condition. A de Sitter horizon is not present as $\Lambda^{\pm}\leq 0$. For 3-D plot, see figure (15).
Figure 4: Schwarzschild – Schwarzschild - de Sitter Junction: Black hole is similar to Schwarzschild - anti-de Sitter – Schwarzschild (figure 3) with a similar asymptote, though in this case the de Sitter Horizon is also present. Again $[C]=0.001$ in this case, fulfilling the asymptote condition. De Sitter Horizon is present.
Figure 5: Schwarzschild – Schwarzschild - anti-de Sitter Junction: Black hole is similar to Schwarzschild - de Sitter – Schwarzschild (figure 2) but does not have a de Sitter Horizon. [C]=0.001[C]=-0.001 and there is no asymptote.
@@
Throughout this work, we define $A]\equiv A^{+}-A^{-}$ and $\bar{A}\equiv\frac{1}{2}(A^{+}+A^{-})$ for some quantity $A$. The first Darmois condition for joining a portion of $\mathcal{M}^{+}$ to a portion of $\mathcal{M}^{-}$ is (see, e.g., [22])
$g_{ij}]=0.$ (1)
@@
Figure 6: Schwarzschild - de Sitter – Schwarzschild - de Sitter Junction: Black hole still possesses asymptote as $[C]=0$ fulfills asymptote condition in equation (27) if there exists a $\Lambda>0$. The location and shape of the asymptote differs from other cases and occurs at much higher $R$. Black hole plot has been expanded to view unstable region at $P^{\prime}/\sigma^{\prime}<-12$.
@@
This boundary causes a discontinuity in the extrinsic curvature (second fundamental form) of the union of $\mathcal{M}^{+}$ and $\mathcal{M}^{-}$. The stress-energy tensor ($S_{ij}$) for this boundary can be calculated using the Lanczos equation, which is given by [22]
@@
We treat $S_{ij}$ analogously to a perfect 4-fluid with $S_{ij}={\rm diag}(-\sigma,P,P)$ [24]. It can be shown that the energy density is
@@
If equation (1) holds and $K_{\theta\theta}]=0$ then we refer to $\Sigma$ as a boundary surface, if $K_{\theta\theta}]\neq 0$ then $\Sigma$ is a thin shell.
00footnotetext: The horizon occurs at $R/m^{+}\approx 54$ which is close to $\alpha_{\rm dS}\approx 54.77$ for $C=0.001$ – see text
In Section III, we consider a junction between two spacetimes of the form,
@@
where $w$ is determined by the direction of the normal vectors. If the vectors point in the opposite directions w=1w=-1 and the junction is referred to as a wormhole. If the vectors point in the same direction $w=1$ and the junction is referred to as a black hole.
@@
For this static solution, $V(R_{0})=V^{\prime}(R_{0})=0$. A stable solution is given by $V^{\prime\prime}(R_{0})>0$. It follows from the definition of $V(R)$ that the equilibrium condition $V^{\prime}(R)=0$ for a static solution is
$\bigg(\frac{M}{2R}\bigg)^{\prime}=-\frac{2R}{M}\bigg(\bigg(\frac{[\mu]}{R}\bigg)\bigg(\frac{[\mu]}{R}\bigg)^{\prime}+\bigg(\frac{\bar{\mu}}{R}\bigg)^{\prime}\bigg)\equiv\Gamma.$ (13)
@@
$\dot{\sigma}=-2\frac{\dot{R}}{R}(\sigma+P)+\Xi,$ (16)
@@
Note that the stability regions lie above and below the surface mapped out by the transparency condition (equation 19).
Stability within the region defined by $0\leq\frac{P^{\prime}}{\sigma^{\prime}}<1$ is of particular interest as it is the causal region. $P^{\prime}/\sigma^{\prime}$ corresponds to the square of a sound speed of perturbations. Thus, for physical solutions it seems natural to restrict this sound speed to real, sub-luminal values, where it is causal. The legitimacy of this assumption is discussed in the conclusion.
@@ III Schwarzschild and (Anti-) de Sitter spacetimes
We begin our stability analysis by considering Schwarzschild wormhole and black hole solutions. We also consider the effects of a positive and negative cosmological constant corresponding to Schwarzschild - de Sitter and Schwarzschild - anti-de Sitter spacetime respectively.
@@
Throughout the rest of this paper where relevant, we only consider cases where $m^{-}<m^{+}$, which is reflected in our axis limits of our plots $0<m^{-}/m^{+}<1$ [8] (in general $\mu^{-}$ and $\mu^{+}$ can be swapped without consequence to the stability regions).
@@
An asymptote between these two horizons occurs if $\Upsilon=0$ [8]. A wormhole with any combination of the above spacetimes always contains such an asymptote. A black hole may or may not contain an asymptote based on the following.
It can be seen from plotting that for small $R$ (near the event horizon) $\Upsilon$ and $M$ both have the same sign. With the exception of Schwarzschild – Schwarzschild (figure 1), after a certain point (before $\Upsilon=0$ or $M=0$) if $\Upsilon$ monotonically increases, $M$ monotonically decreases or vice versa. Thus, only one of these quantities can be equal to zero for a certain junction. If $M=0$ anywhere then $\Upsilon\neq 0$ and there is no asymptote, though a stability flip would still exist at $M=0$. This lets the simpler condition $M=0$ become an indicator of the nonexistence of an asymptote.
@@
Since $R$ must be positive, if $[m]$ and $\Lambda]$ have the same sign, the condition can never be satisfied for any $R$, and so an asymptote must exist. This gives an asymptote existence condition of
@@
For a junction with at least one $\Lambda>0$ there is an upper limit to $R$. If the $R$ value which satisfies the horizon condition (equation 25) is greater than $R_{\rm dS}$ then the condition for asymptote non-existence cannot be met and an asymptote would be present. Taking $R_{\rm dS}\approx\sqrt{3/\Lambda_{\rm max}}$ gives an asymptote existence condition of
@@
For the Schwarzschild – Schwarzschild case (figure 1), $M$ and $\Upsilon$ approach zero for large $R$, but never reach it. Thus, there is neither an asymptote or stability flip in this simple case.
It is also apparent from plotting that $M\Upsilon$ monotonically decreases for black holes and monotonically increases for wormholes. This behavior dictates the location of stability regions, whether above or below the surface.
Figure 7: Schwarzschild - anti-de Sitter – Schwarzschild - anti-de Sitter Junction: For the black hole, $M\neq 0$ for any $\beta$ and there is no stability flip. For the wormhole, the stability region approaches the lower bound of the Causal Region as $\alpha\rightarrow\infty$ but never reaches it.
@@ IV FLRW spacetime
Unlike the previous stationary spacetimes, the FLRW spacetime is expanding. Here we follow the precedent established in the Swiss Cheese Cosmological Model, which considers spherical regions of Schwarzschild spacetime matched to an FLRW background [3]. In these cases, it is typical to define a hypersurface that is expanding at the same rate as the FLRW so that in the FLRW frame it is stationary apart from the evolution of the hypersurface radius defined in equation (11). The surface in the FLRW frame is defined by $r_{\rm f}=R(\tau)$. The static solution solved for here is the one for which $\dot{R}=0$.
@@
$\Xi=\pm\Bigg[\dot{r}(\rho+p){\bigg(\bigg|g^{\alpha\beta}\frac{\partial F}{\partial x^{\alpha}}\frac{\partial F}{\partial x^{\beta}}\bigg|\bigg)^{-1/2}}\Bigg],$ (30)
@@
Here we let $M=-[K_{\theta\theta}]=4\pi\chi^{2}\sigma$. Using this, the continuity equation can finally be expressed as
@@
Figure 8: FLRW (k=0) – Schwarzschild Junction: The black hole and wormhole graphs are similar to previous examples though does not posses a de Sitter Horizon, despite there existing a $\Lambda>0$. It is worth noting that the radius of the event horizon decreases below $\alpha=2$ at low $a$ due to a high value of $H^{2}$. For 3-D plot, see figure (16).
Figure 9: FLRW (k=0) – Schwarzschild - de Sitter Junction: Very similar to figure (8). The $M=0$ line occurs at higher $\alpha$.
@@
For FLRW, $K_{\theta\theta}^{-}$ yields a complicated form. We simplify this by taking $K_{\theta\theta}$ at the static solution $\dot{R}=\frac{dR}{dt}\frac{dt}{d\tau}=0\Rightarrow\frac{dR}{dt}=0$ if $\frac{dt}{d\tau}\neq 0$.
@@
Though, at the static solution $V=H\chi\neq 0$ generally. $V^{\prime}$ may be be chosen to be 0 for an equilibrium point. Thus, $V^{\prime\prime}>0$ again yields stable equilibria. Combining this with the transparency condition (equation 35) and the fact that $V^{\prime}=0$ at this point, gives stability conditions of
@@
Figure 10: FLRW (k=0) – Schwarzschild - anti-de Sitter Junction: Very similar to figure (8). The $M=0$ line occurs at lower $\alpha$.
@@
Figure 11: FLRW (k=+1) – Schwarzschild Junction: For $k\neq 0$ we choose $m=0.02$. It is important to note that $k>0$ gives the non physical result of $H^{2}<0$ for $0.420\lesssim a(t)\lesssim 0.823$. A horizon caused by the positive curvature exists at high $\chi$ for both black hole and wormhole. At larger $m$ values the asymptote condition can be fulfilled but commonly this occurs only when $H^{2}<0$. For 3-D plot, see figure (17).